# An Adaptive Approach for Infinitely Many-armed Bandits under Generalized Rotting Constraints

### v3

@@ Conclusion:
Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$ are i.i.d. random variables with geometric distribution with parameter $2\delta$, given a fixed value of $m^{\mathcal{G}}$. Therefore, for non-negative integer $k$ we have $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-2\delta)^{k}2\delta$, for $i=1,\ldots,m^{\mathcal{G}}$.
@@
For $2\delta<x\leq 1$, let $b(x)=\mathbb{P}(\Delta_{t_{1}}(a)=x|a\text{ is a bad arm})$. Then we have $\mathbb{P}(\Delta_{t_{1}}(a)=x|a\text{ is a bad arm})=\mathbb{P}(\Delta_{t_{1}}(a)=x|\Delta_{t_{1}}(a)>2\delta)=\mathbb{P}(\Delta_{t_{1}}(a)=x)/\mathbb{P}(\Delta_{1}(a)>2\delta)=\mathbb{P}(\Delta_{t_{1}}(a)=x)/(1-C(2\delta)^{\beta})=O(\mathbb{P}(\Delta_{t_{1}}(a)=x))$, where the last equality comes from small $\delta$ with $S_{T}=o(T)$. For any $i\in[m^{\mathcal{G}}]$, $j\in[m^{\mathcal{B}}_{i}]$, we have
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta=\max\{(S_{T}/T)^{1/(\beta+1)},(S_{T}/T)^{1/2}\}$ and $m^{\mathcal{G}}=S_{T}$, from Lemma  A.7 and the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables following geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=1/(2\delta)^{\beta}-1$, we have
@@ A.6 Details for the Case of Unknown Parameters
We define $\delta^{\dagger}_{V}=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3},1/H^{1/(\beta+1)},1/\sqrt{H}\}$ and $\delta^{\dagger}_{S}=\max\{(S_{T}/T)^{1/(\beta+1)},(S_{T}/T)^{1/2},1/H^{1/(\beta+1)},1/\sqrt{H}\}$. Then the optimized threshold parameter is $\delta^{\dagger}_{VS}=\min\{\delta_{S}^{\dagger},\delta_{V}^{\dagger}\}$. The optimized threshold parameter can be derived from the theoretical analysis in Appendix A.7. The target of the master is to find the parameter. From the above, we can observe that $1/\sqrt{H}\leq\delta^{\dagger}_{VS}\leq 1$. Therefore, we set =$\mathcal{B}=\{1/2,\dots,1/2^{\log_{2}\sqrt{H}}\}$ which is the candidate values for unknown $\delta^{\dagger}$.
@@
We can observe that there is the additional regret cost of $T^{(2\beta+1)/(2\beta+2)}$ for $\beta\geq 1$ or $T^{3/4}$ for $10<\beta<$ 1 compared to Algorithm 1. This additional cost originates from the additional procedure to learn the optimal value of $\delta$ in Algorithm 2, which is negligible when $V_{T}$ and $S_{T}$ are large enough. We also note that the well-known black-box framework proposed for addressing nonstationarity [25] is not applicable to this problem because the UCB of the chosen arm in this context does not consistently surpass the optimal mean reward, violating the necessary assumption for the framework. Attaining the optimal regret bound under a parameter-free algorithm remains an unresolved issue.
@@ A.7.1 Case of $\delta_{V}^{\dagger}\leq\delta_{S}^{\dagger}$
Note that $R_{1}^{\pi}(T)$ accounts for the regret caused by the near-optimal base algorithm $\pi_{i}(\delta^{\dagger})$’s against the optimal mean reward and $R_{2}^{\pi}(T)$ accounts for the regret caused by the master algorithm by selecting a base with $\delta\in\mathcal{B}$ at every block against the base with $\delta^{\dagger}$. In what follows, we provide upper bounds for each regret component. We first provide an upper bound for $\mathbb{E}[R_{1}^{\pi}(T)]$ by following the proof steps in Theorem 3.1. Then we provide an upper bound for $\mathbb{E}[R_{2}^{\pi}(T)]$. We set $H=\lceil T^{1/2}\rceil$ and $\delta^{\dagger}$ to be a smallest value in $\mathcal{B}$ which is larger than $\delta^{\dagger}_{V}=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3},1/H^{1/(\beta+1)},1/H^{1/2}\}$.
@@
Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$ are i.i.d. random variables with geometric distribution with parameter $2\delta$, given a fixed value of $m^{\mathcal{G}}$. Therefore, for non-negative integer $k$ we have $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-2\delta^{\dagger})^{k}2\delta^{\dagger}$, for $i=1,\ldots,m^{\mathcal{G}}$. Define $\tilde{m}_{H}$ to be the number of episodes from the policy $\pi$ over the horizon $H$, $\tilde{m}_{H}^{\mathcal{G}}$ to be the total number of samples of a good arm by the policy $\pi$ over the horizon $H$ such that $\tilde{m}_{H}^{\mathcal{G}}=\tilde{m}_{H}$ or $\tilde{m}_{H}^{\mathcal{G}}=\tilde{m}-1$, and $\tilde{m}_{i,H}^{\mathcal{B}}$ to be the number of samples of a bad arm in the $i$-th episode by the policy $\pi_{1}(\delta^{\dagger})$ over the horizon $H$.
@@
We first consider the case where $V_{T}>\max\{T/H^{3/2},T/H^{(\beta+2)/(\beta+1)}\}$. In this case, we have $\delta^{\dagger}=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$. Here, we define the policy $\pi$ after time $H$ such that it pulls a good arm until its total rotting variation is equal to or greater than $\delta^{\dagger}/2$ and does not pull a sampled bad arm. We note that defining how $\pi$ works after $H$ is only for the proof to get a regret bound over time horizon $H$. For the last arm $\tilde{a}$ over the horizon $H$, it pulls the arm until its total variation becomes $\max\{\delta^{\dagger}/2,V_{H}(\tilde{a})\}$ if $\tilde{a}$ is a good arm. For $i\in[m^{\mathcal{G}}]$, $j\in[m_{i}^{\mathcal{B}}]$ let $V_{i}^{\mathcal{G}}$ and $V_{i,j}^{\mathcal{B}}$ be the total rotting variation of pulling the good arm in $i$-th episode and $j$-th bad arm in $i$-th episode from the policy, respectively. Here we define $V_{i}^{\mathcal{G}}$’s and $V_{i,j}^{\mathcal{B}}$’s as follows:
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta^{\dagger}=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$ and $m^{\mathcal{G}}=\lceil 2V_{H,1}/\delta^{\dagger}\rceil$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=1/(2\delta^{\dagger})^{\beta}-1$, we have
@@
From $R^{\pi_{1}(\delta^{\dagger})}(H)=R^{\mathcal{G}}(H)+R^{\mathcal{B}}(H)$ and Lemmas A.17, A.19, A.20, with $\delta^{\dagger}=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$ we have
@@
Now, we consider the case where $V_{T}\leq\max\{T/H^{3/2},T/H^{(\beta+2)/(\beta+1)}\}$. In this case, we have $\delta^{\dagger}=\max\{1/\sqrt{H},1/H^{\frac{1}{\beta+1}}\}$. From the result of Lemma A.19, by setting $m^{\mathcal{G}}=\lceil 2V_{H,1}/\delta^{\dagger}\rceil$ we have $R^{\mathcal{B}}(H)\leq R^{{\mathcal{B}}}_{m^{\mathcal{G}}}$.
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta^{\dagger}=\max\{1/H^{1/2},1/H^{1/(\beta+1)}\}$ and $m^{\mathcal{G}}=\lceil 2V_{H,1}/\delta^{\dagger}\rceil$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=1/(2\delta^{\dagger})^{\beta}-1$, we have
@@
From $R^{\pi_{1}(\delta^{\dagger})}(H)=R^{\mathcal{G}}(H)+R^{\mathcal{B}}(H)$ and Lemmas A.17, A.19, A.21, with $\delta^{\dagger}=\Theta(\max\{1/H^{1/2},1/H^{1/(\beta+1)}\})$ we have
@@ A.7.2 Case of $\delta_{V}^{\dagger}>\delta_{S}^{\dagger}$
Note that $R_{1}^{\pi}(T)$ accounts for the regret caused by the near-optimal base algorithm $\pi_{i}(\delta^{\dagger})$’s against the optimal mean reward and $R_{2}^{\pi}(T)$ accounts for the regret caused by the master algorithm by selecting a base with $\delta\in\mathcal{B}$ at every block against the base with $\delta^{\dagger}$. In what follows, we provide upper bounds for each regret component. We first provide an upper bound for $\mathbb{E}[R_{1}^{\pi}(T)]$ by following the proof steps in Theorem 3.3. Then we provide an upper bound for $\mathbb{E}[R_{2}^{\pi}(T)]$. We set $\delta^{\dagger}$ to be a smallest value in $\mathcal{B}$ which is larger than $\delta_{S}^{\dagger}=\max\{(S_{T}/T)^{1/(\beta+1)},1/H^{1/(\beta+1)},(S_{T}/T)^{1/2},1/H^{1/2}\}$ such that we have $\delta^{\dagger}=\Theta(\max\{(S_{T}/T)^{1/(\beta+1)},1/H^{1/(\beta+1)},(S_{T}/T)^{1/2},1/H^{1/2}\})$.
@@
Now we analyze $R^{\mathcal{B}}(H)$. We divide regret $R^{\mathcal{B}}(H)$ into two regret from bad arms in $\mathcal{A}(1)/\mathcal{A}_{S}(1)$, denoted by $R^{{\mathcal{B}},1}(H)$, and regret from bad arms in $\mathcal{A}_{S}(1)$, denoted by $R^{{\mathcal{B}},2}(H)$ such that $R^{\mathcal{B}}(H)=R^{{\mathcal{B}},1}(H)+R^{{\mathcal{B}},2}(H)$. We first analyze $R^{{\mathcal{B}},1}(H)$ in the following. We consider arms in $\mathcal{A}(1)/\mathcal{A}_{S}(1)$. For the proof, we adopt the episodic approach in Kim et al. [13] for regret analysis. In the following, we introduce some notation. Here we only consider arms in $\mathcal{A}(1)/\mathcal{A}_{S}(1)$ so that the following notation is defined without considering (rotted) arms in $\mathcal{A}_{S}(1)$. Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$’s are i.i.d. random variables with geometric distribution with parameter $2\delta^{\dagger}$, conditional on the value of $m^{\mathcal{G}}$. Therefore, $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-2\delta^{\dagger})^{k}2\delta^{\dagger}$, for $i=1,\ldots,m^{\mathcal{G}}$.
@@
From Lemmas A.23, A.24, A.25, we have

### v4

@@ Conclusion:
Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$ are i.i.d. random variables with geometric distribution with parameter $C(2\delta)^{\beta}$ for some constant $C>0$, given a fixed value of $m^{\mathcal{G}}$. Therefore, for non-negative integer $k$ we have $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-C(2\delta)^{\beta})^{k}C(2\delta)^{\beta}$, for $i=1,\ldots,m^{\mathcal{G}}$.
@@
For $2\delta<x\leq 1$, let $b(x)=\mathbb{P}(\Delta_{t_{1}}(a)=x|a\text{ is a bad arm})$. Then we have $\mathbb{P}(\Delta_{t_{1}}(a)=x|a\text{ is a bad arm})=\mathbb{P}(\Delta_{t_{1}}(a)=x|\Delta_{t_{1}}(a)>2\delta)=\mathbb{P}(\Delta_{t_{1}}(a)=x)/\mathbb{P}(\Delta_{1}(a)>2\delta)=\mathbb{P}(\Delta_{t_{1}}(a)=x)/(1-C(2\delta)^{\beta})=O(\mathbb{P}(\Delta_{t_{1}}(a)=x))$, where the last equality comes from small $\delta$ with small enough $c_{1}<1$. For any $i\in[m^{\mathcal{G}}]$, $j\in[m^{\mathcal{B}}_{i}]$, we have
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta=c_{1}\max\{(S_{T}/T)^{1/(\beta+1)},(S_{T}/T)^{1/2}\}$ and $m^{\mathcal{G}}=S_{T}$, from Lemma  A.7 and the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables following geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=(1/C(2\delta)^{\beta})-1$ for some constant $C>0$, we have
@@ A.6 Details for the Case of Unknown Parameters
We define $\delta^{\dagger}_{V}=c_{1}\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3},1/H^{1/(\beta+1)},1/\sqrt{H}\}$ and $\delta^{\dagger}_{S}=c_{1}\max\{(S_{T}/T)^{1/(\beta+1)},(S_{T}/T)^{1/2},1/H^{1/(\beta+1)},1/\sqrt{H}\}$ for some constant $10<c_{1}<$ 1. Then the optimized threshold parameter is $\delta^{\dagger}_{VS}=\min\{\delta_{S}^{\dagger},\delta_{V}^{\dagger}\}$. The optimized threshold parameter can be derived from the theoretical analysis in Appendix A.7. The target of the master is to find the parameter. From the above, we can observe that $c_{1}/\sqrt{H}\leq\delta^{\dagger}_{VS}\leq 1$. Therefore, we set =$\mathcal{B}=\{1/2,\dots,1/2^{\log_{2}\sqrt{H}/c_{1}}\}$ which is the candidate values for unknown $\delta^{\dagger}$.
@@
We can observe that there is the additional regret cost of $T^{(2\beta+1)/(2\beta+2)}$ for $\beta\geq 1$ or $T^{3/4}$ for $10<\beta<$ 1 compared to Algorithm 1. This additional cost originates from the additional procedure to learn the optimal value of $\delta$ in Algorithm 2, which is negligible when $V_{T}$ and $S_{T}$ are large enough.
Remark A.16.
In the case where the value of $\beta$ is known, setting $H=\lceil\max\{T^{(\beta+1)/(\beta+3)},\sqrt{T}\}\rceil$ reduces the additional regret cost of Algorithm 2 to $\max\{T^{(\beta+2)/(\beta+3)},T^{3/4}\}$.
Attaining the optimal regret bound under a parameter-free algorithm remains an open problem.
@@ A.7.1 Case of $\delta_{V}^{\dagger}\leq\delta_{S}^{\dagger}$
Note that $R_{1}^{\pi}(T)$ accounts for the regret caused by the near-optimal base algorithm $\pi_{i}(\delta^{\dagger})$’s against the optimal mean reward and $R_{2}^{\pi}(T)$ accounts for the regret caused by the master algorithm by selecting a base with $\delta\in\mathcal{B}$ at every block against the base with $\delta^{\dagger}$. In what follows, we provide upper bounds for each regret component. We first provide an upper bound for $\mathbb{E}[R_{1}^{\pi}(T)]$ by following the proof steps in Theorem 3.1. Then we provide an upper bound for $\mathbb{E}[R_{2}^{\pi}(T)]$. We set $H=\lceil T^{1/2}\rceil$ and $\delta^{\dagger}$ to be a smallest value in $\mathcal{B}$ which is larger than $\delta^{\dagger}_{V}=c_{1}\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3},1/H^{1/(\beta+1)},1/H^{1/2}\}$.
@@
Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$ are i.i.d. random variables with geometric distribution with parameter $2\delta$, given a fixed value of $m^{\mathcal{G}}$. Therefore, for non-negative integer $k$ we have $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-C(2\delta^{\dagger})^{\beta})^{k}C(2\delta^{\dagger})^{\beta}$ for some constant $C>0$, for $i=1,\ldots,m^{\mathcal{G}}$. Define $\tilde{m}_{H}$ to be the number of episodes from the policy $\pi$ over the horizon $H$, $\tilde{m}_{H}^{\mathcal{G}}$ to be the total number of samples of a good arm by the policy $\pi$ over the horizon $H$ such that $\tilde{m}_{H}^{\mathcal{G}}=\tilde{m}_{H}$ or $\tilde{m}_{H}^{\mathcal{G}}=\tilde{m}-1$, and $\tilde{m}_{i,H}^{\mathcal{B}}$ to be the number of samples of a bad arm in the $i$-th episode by the policy $\pi_{1}(\delta^{\dagger})$ over the horizon $H$.
@@
We first consider the case where $V_{T}>\max\{T/H^{3/2},T/H^{(\beta+2)/(\beta+1)}\}$. In this case, we have $\delta^{\dagger}=c_{1}\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$. Here, we define the policy $\pi$ after time $H$ such that it pulls a good arm until its total rotting variation is equal to or greater than $\delta^{\dagger}/2$ and does not pull a sampled bad arm. We note that defining how $\pi$ works after $H$ is only for the proof to get a regret bound over time horizon $H$. For the last arm $\tilde{a}$ over the horizon $H$, it pulls the arm until its total variation becomes $\max\{\delta^{\dagger}/2,V_{H}(\tilde{a})\}$ if $\tilde{a}$ is a good arm. For $i\in[m^{\mathcal{G}}]$, $j\in[m_{i}^{\mathcal{B}}]$ let $V_{i}^{\mathcal{G}}$ and $V_{i,j}^{\mathcal{B}}$ be the total rotting variation of pulling the good arm in $i$-th episode and $j$-th bad arm in $i$-th episode from the policy, respectively. Here we define $V_{i}^{\mathcal{G}}$’s and $V_{i,j}^{\mathcal{B}}$’s as follows:
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta^{\dagger}=c_{1}\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$ and $m^{\mathcal{G}}=\lceil 2V_{H,1}/\delta^{\dagger}\rceil$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=1/(2\delta^{\dagger})^{\beta}-1$, we have
@@
From $R^{\pi_{1}(\delta^{\dagger})}(H)=R^{\mathcal{G}}(H)+R^{\mathcal{B}}(H)$ and Lemmas A.18, A.20, A.21, with $\delta^{\dagger}=\max\{(V_{T}/T)^{1/(\beta+2)},(V_{T}/T)^{1/3}\}$ we have
@@
Now, we consider the case where $V_{T}\leq\max\{T/H^{3/2},T/H^{(\beta+2)/(\beta+1)}\}$. In this case, we have $\delta^{\dagger}=c_{1}\max\{1/\sqrt{H},1/H^{\frac{1}{\beta+1}}\}$. From the result of Lemma A.20, by setting $m^{\mathcal{G}}=\lceil 2V_{H,1}/\delta^{\dagger}\rceil$ we have $R^{\mathcal{B}}(H)\leq R^{{\mathcal{B}}}_{m^{\mathcal{G}}}$.
@@
Recall that $R^{{\mathcal{B}}}_{m^{\mathcal{G}}}=\sum_{i=1}^{m^{\mathcal{G}}}\sum_{j\in[m_{i}^{\mathcal{B}}]}R_{i,j}^{\mathcal{B}}.$ With $\delta^{\dagger}=c_{1}\max\{1/H^{1/2},1/H^{1/(\beta+1)}\}$ and $m^{\mathcal{G}}=\lceil 2V_{H,1}/\delta^{\dagger}\rceil$, from the fact that $m_{i}^{\mathcal{B}}$’s are i.i.d. random variables with geometric distribution with $\mathbb{E}[m_{i}^{\mathcal{B}}]=(1/C(2\delta^{\dagger})^{\beta})-1$, we have
@@
From $R^{\pi_{1}(\delta^{\dagger})}(H)=R^{\mathcal{G}}(H)+R^{\mathcal{B}}(H)$ and Lemmas A.18, A.20, A.22, with $\delta^{\dagger}=\Theta(\max\{1/H^{1/2},1/H^{1/(\beta+1)}\})$ we have
@@ A.7.2 Case of $\delta_{V}^{\dagger}>\delta_{S}^{\dagger}$
Note that $R_{1}^{\pi}(T)$ accounts for the regret caused by the near-optimal base algorithm $\pi_{i}(\delta^{\dagger})$’s against the optimal mean reward and $R_{2}^{\pi}(T)$ accounts for the regret caused by the master algorithm by selecting a base with $\delta\in\mathcal{B}$ at every block against the base with $\delta^{\dagger}$. In what follows, we provide upper bounds for each regret component. We first provide an upper bound for $\mathbb{E}[R_{1}^{\pi}(T)]$ by following the proof steps in Theorem 3.3. Then we provide an upper bound for $\mathbb{E}[R_{2}^{\pi}(T)]$. We set $\delta^{\dagger}$ to be a smallest value in $\mathcal{B}$ which is larger than $\delta_{S}^{\dagger}=c_{1}\max\{(S_{T}/T)^{1/(\beta+1)},1/H^{1/(\beta+1)},(S_{T}/T)^{1/2},1/H^{1/2}\}$ such that we have $\delta^{\dagger}=\Theta(\max\{(S_{T}/T)^{1/(\beta+1)},1/H^{1/(\beta+1)},(S_{T}/T)^{1/2},1/H^{1/2}\})$.
@@
Now we analyze $R^{\mathcal{B}}(H)$. We divide regret $R^{\mathcal{B}}(H)$ into two regret from bad arms in $\mathcal{A}(1)/\mathcal{A}_{S}(1)$, denoted by $R^{{\mathcal{B}},1}(H)$, and regret from bad arms in $\mathcal{A}_{S}(1)$, denoted by $R^{{\mathcal{B}},2}(H)$ such that $R^{\mathcal{B}}(H)=R^{{\mathcal{B}},1}(H)+R^{{\mathcal{B}},2}(H)$. We first analyze $R^{{\mathcal{B}},1}(H)$ in the following. We consider arms in $\mathcal{A}(1)/\mathcal{A}_{S}(1)$. For the proof, we adopt the episodic approach in Kim et al. [13] for regret analysis. In the following, we introduce some notation. Here we only consider arms in $\mathcal{A}(1)/\mathcal{A}_{S}(1)$ so that the following notation is defined without considering (rotted) arms in $\mathcal{A}_{S}(1)$. Given a policy sampling arms in the sequence order, let $m^{\mathcal{G}}$ be the number of samples of distinct good arms and $m^{{\mathcal{B}}}_{i}$ be the number of consecutive samples of distinct bad arms between the $i-1$-st and $i$-th sample of a good arm among $m^{\mathcal{G}}$ good arms. We refer to the period starting from sampling the $i-1$-st good arm before sampling the $i$-th good arm as the $i$-th episode. Observe that $m^{\mathcal{B}}_{1},\ldots,m^{\mathcal{B}}_{m^{\mathcal{G}}}$’s are i.i.d. random variables with geometric distribution with parameter $C(2\delta^{\dagger})^{\beta}$ for some constant $C>0$, conditional on the value of $m^{\mathcal{G}}$. Therefore, $\mathbb{P}(m^{\mathcal{B}}_{i}=k)=(1-C(2\delta^{\dagger})^{\beta})^{k}C(2\delta^{\dagger})^{\beta}$, for $i=1,\ldots,m^{\mathcal{G}}$.
@@
From Lemmas A.24, A.25, A.26, we have